# labelinglabel

### v2

@@ 3. Generalized wreath products
Definition 3.3.
Let $G$ be a countable discrete group, $A$ be a finite set. Let $\alpha_{A}:\textrm{Sym}(A)\curvearrowright A$ be the natural action. We may then consider the generalized wreath product $G\wr_{\alpha_{A}}\textrm{Sym}(A)$, which is of the form $G^{\oplus A}\rtimes_{\beta_{A}}\textrm{Sym}(A)$ where $\beta_{A}$ is the action $\beta_{A}:\textrm{Sym}(A)\curvearrowright G^{\oplus A}$ induced by $\alpha_{A}$. We define a metric $d_{G,A}$ on this group given by
@@
where $g_{i}\in G^{\oplus A}$ and $\sigma_{i}\in\textrm{Sym}(A)$, with the former regarded as functions from $A$ to $G$. It is easy to verify that this is indeed a metric and is invariant under both left and right multiplication.
@@
[…]
@@
[…]
@@
$\begin{split}P^{i}_{s}(\beta_{h_{1}}(g_{2}))(b)&=\begin{cases}g_{2}(\alpha(h_{1}^{-1})x),\textrm{if }x\in\alpha(h_{1})\cdot\textrm{supp}(g_{2})\textrm{ and }\pi^{i}_{s}(x)=b\\ 1_{G},\textrm{otherwise}\end{cases}\\ &=\begin{cases}g_{2}(x),\textrm{if }x\in\textrm{supp}(g_{2})\textrm{ and }\pi^{i}_{s}(\alpha(h_{1})x)=b\\ 1_{G},\textrm{otherwise}\end{cases}\\ &=\begin{cases}g_{2}(x),\textrm{if }x\in\textrm{supp}(g_{2})\textrm{ and }\pi^{i}_{\varphi_{i}(h_{1}^{-1})s}(x)=b\\ 1_{G},\textrm{otherwise}\end{cases}\\ &=\begin{cases}g_{2}(x),\textrm{if }x\in E_{i}\textrm{ and }\pi^{i}_{\varphi_{i}(h_{1}^{-1})s}(x)=b\\ 1_{G},\textrm{otherwise}\end{cases}\\ &=P^{i}_{\varphi_{i}(h_{1}^{-1})s}(g_{2})(b)\end{split}$
@@
Let $H$ be a sofic group, $\alpha:H\curvearrowright X$ be a sofic action, $A\leq G$ be an inclusion of countable discrete groups s.t. the amalgamated free product of any countably many copies of $G$ over $A$ is sofic. Then the amalgamated free generalized wreath product $G\wr^{\ast_{A}}_{\alpha}H$ is sofic. In particular, if $G,H$ are sofic groups and $\alpha:H\curvearrowright X$ is a sofic action, then the free generalized wreath product $G\wr^{\ast}_{\alpha}H$ is sofic, and under the same conditions, if we in addition have an amenable subgroup $A$ of $G$, then the amalgamated free generalized wreath product $G\wr^{\ast_{A}}_{\alpha}H$ is sofic.
@@
Again, fix increasing sequences of finite subsets $F_{1}\subseteq F_{2}\subseteq\cdots\subseteq H$ and $E_{1}\subseteq E_{2}\subseteq\cdots\subseteq X$ s.t. $\cup_{i}F_{i}=H$, $\cup_{i}E_{i}=X$. Fix a decreasing sequence $\epsilon_{i}>0$ s.t. $\lim_{i}\epsilon_{i}=0$. For each $i$, let $\varphi_{i}:H\rightarrow\textrm{Sym}(A_{i})$ be a unital, $(F_{i},\epsilon_{i})$-multiplicative, and an $(F_{i},E_{i},\epsilon_{i})$-orbit approximation of $\alpha$. By definition of $(F_{i},E_{i},\epsilon_{i})$-orbit approximation, there exists a finite set $B_{i}$ and a subset $S_{i}\subseteq A_{i}$ s.t. $|S_{i}|>(1-\epsilon_{i})|A_{i}|$ and for each $s\in S_{i}$ there is an injective map $\pi^{i}_{s}:E_{i}\hookrightarrow B_{i}$ s.t. $\pi^{i}_{\varphi_{i}(g)s}(x)=\pi^{i}_{s}(\alpha(g^{-1})x)$ for all $s\in S_{i}$, $g\in F_{i}$, $x\in E_{i}$, whenever $\varphi_{i}(g)s\in S_{i}$ and $\alpha(g^{-1})x\in E_{i}$. Let $G_{i}=\ast^{b\in B_{i}}_{A}G_{b}$, where $G_{b}$ are copies of $G$. Under the assumption that the amalgamated free product of any countably many copies of $G$ over $A$ is sofic, we see that $G_{i}$ is sofic. Let $p_{i}:\ast^{x\in X}_{A}G_{x}\rightarrow\ast^{x\in E_{i}}_{A}G_{x}$, where $G_{x}$ are copies of $G$, be the map that is the identity map on $\ast^{x\in E_{i}}_{A}G_{x}$ and sends everything else to $1$. Let $q^{i}_{s}:\ast^{x\in E_{i}}_{A}G_{x}\hookrightarrow\ast^{b\in B_{i}}_{A}G_{b}=G_{i}$ be the inclusion map induced by $\pi^{i}_{s}:E_{i}\hookrightarrow B_{i}$. Let $P^{i}_{s}=q^{i}_{s}\circ p_{i}$.
@@
Let $H$ be a hyperlinear group, $\alpha:H\curvearrowright X$ be a sofic action, $N\subseteq M$ be an inclusion of tracial von Neumann algebras s.t. the amalgamated free product of any countably many copies of $M$ over $N$ is Connes-embeddable. Then the amalgamated free generalized wreath product $M\wr^{\ast_{N}}_{\alpha}H$ is Connes-embeddable. In particular, if $H$ is a hyperlinear group, $\alpha:H\curvearrowright X$ is a sofic action, and $M$ is a Connes-embeddable tracial von Neumann algebra, then the free generalized wreath product $M\wr^{\ast}_{\alpha}H$ is Connes-embeddable, and under the same conditions, if we in addition have an amenable subalgebra $N$ of $M$, then the amalgamated free generalized wreath product $M\wr^{\ast_{N}}_{\alpha}H$ is Connes-embeddable.
@@
Let $H$ be a hyperlinear group, $\alpha:H\curvearrowright X$ be a sofic action, $A\leq G$ be an inclusion of countable discrete groups s.t. the amalgamated free product of any countably many copies of $G$ over $A$ is hyperlinear. Then the amalgamated free generalized wreath product $G\wr^{\ast_{A}}_{\alpha}H$ is hyperlinear. In particular, if $G,H$ are hyperlinear groups and $\alpha:H\curvearrowright X$ is a sofic action, then the free generalized wreath product $G\wr^{\ast}_{\alpha}H$ is hyperlinear, and under the same conditions, if we in addition have an amenable subgroup $A$ of $G$, then the amalgamated free generalized wreath product $G\wr^{\ast_{A}}_{\alpha}H$ is hyperlinear.

### v3

@@ 3. Generalized wreath products
Definition $3.$
Let $G$ be a countable discrete group, AA be a finite set. Let $\alpha_{A}:\textrm{Sym}(A)\curvearrowright A$ be the natural action. We may then consider the generalized wreath product $G\wr_{\alpha_{A}}\textrm{Sym}(A)$, which is of the form $G^{\oplus A}\rtimes_{\beta_{A}}\textrm{Sym}(A)$ where $\beta_{A}$ is the action $\beta_{A}:\textrm{Sym}(A)\curvearrowright G^{\oplus A}$ induced by $\alpha_{A}$. We define a metric $d_{G,A}$ on this group given by
@@
where $g_{i}\in G^{\oplus A}$ and $\sigma_{i}\in\textrm{Sym}(A)$, with the former regarded as functions from AA to $G$. It is easy to verify that this is indeed a metric and is invariant under both left and right multiplication.
@@
Let $G$ be a sofic group, AA be a finite set. Then there exists a free ultrafilter $\mathcal{U}$ on $\mathbb{N}$ and a sequence of finite sets $F_{i}$, s.t. $(G^{\oplus A}\rtimes_{\beta_{A}}\textrm{Sym}(A),d_{G,A})$ embeds into $\prod_{\mathcal{U}}(\textrm{Sym}(F_{i}),d)$ isometrically.
@@
Let $G$ be a countable discrete group. Suppose there exists a free ultrafilter $\mathcal{U}$ on $\mathbb{N}$, a sequence of sofic groups $G_{i}$, and a sequence of finite sets $A_{i}$ s.t. $G$ embeds into $\prod_{\mathcal{U}}(G_{i}^{\oplus A_{i}}\rtimes_{\beta_{A_{i}}}\textrm{Sym}(A_{i}),d_{G_{i},A_{i}})$, then $G$ is sofic.
@@
[…]
@@
Let $H$ be a sofic group, $\alpha:H\curvearrowright X$ be a sofic action, $A\leq G$ be an inclusion of countable discrete groups s.t. the amalgamated free product of any countably many copies of $G$ over AA is sofic. Then the amalgamated free generalized wreath product $G\wr^{\ast_{A}}_{\alpha}H$ is sofic. In particular, if $G,H$ are sofic groups and $\alpha:H\curvearrowright X$ is a sofic action, then the free generalized wreath product $G\wr^{\ast}_{\alpha}H$ is sofic, and under the same conditions, if we in addition have an amenable subgroup AA of $G$, then the amalgamated free generalized wreath product $G\wr^{\ast_{A}}_{\alpha}H$ is sofic.
@@
Again, fix increasing sequences of finite subsets $F_{1}\subseteq F_{2}\subseteq\cdots\subseteq H$ and $E_{1}\subseteq E_{2}\subseteq\cdots\subseteq X$ s.t. $\cup_{i}F_{i}=H$, $\cup_{i}E_{i}=X$. Fix a decreasing sequence $\epsilon_{i}>0$ s.t. $\lim_{i}\epsilon_{i}=0$. For each $i$, let $\varphi_{i}:H\rightarrow\textrm{Sym}(A_{i})$ be a unital, $(F_{i},\epsilon_{i})$-multiplicative, and an $(F_{i},E_{i},\epsilon_{i})$-orbit approximation of $\alpha$. By definition of $(F_{i},E_{i},\epsilon_{i})$-orbit approximation, there exists a finite set $B_{i}$ and a subset $S_{i}\subseteq A_{i}$ s.t. $|S_{i}|>(1-\epsilon_{i})|A_{i}|$ and for each $s\in S_{i}$ there is an injective map $\pi^{i}_{s}:E_{i}\hookrightarrow B_{i}$ s.t. $\pi^{i}_{\varphi_{i}(g)s}(x)=\pi^{i}_{s}(\alpha(g^{-1})x)$ for all $s\in S_{i}$, $g\in F_{i}$, $x\in E_{i}$, whenever $\varphi_{i}(g)s\in S_{i}$ and $\alpha(g^{-1})x\in E_{i}$. Let $G_{i}=\ast^{b\in B_{i}}_{A}G_{b}$, where $G_{b}$ are copies of $G$. Under the assumption that the amalgamated free product of any countably many copies of $G$ over AA is sofic, we see that $G_{i}$ is sofic. Let $p_{i}:\ast^{x\in X}_{A}G_{x}\rightarrow\ast^{x\in E_{i}}_{A}G_{x}$, where $G_{x}$ are copies of $G$, be the map that is the identity map on $\ast^{x\in E_{i}}_{A}G_{x}$ and sends everything else to 11. Let $q^{i}_{s}:\ast^{x\in E_{i}}_{A}G_{x}\hookrightarrow\ast^{b\in B_{i}}_{A}G_{b}=G_{i}$ be the inclusion map induced by $\pi^{i}_{s}:E_{i}\hookrightarrow B_{i}$. Let $P^{i}_{s}=q^{i}_{s}\circ p_{i}$.
@@
Let $H$ be a hyperlinear group, $\alpha:H\curvearrowright X$ be a sofic action, $N\subseteq M$ be an inclusion of tracial von Neumann algebras s.t. the amalgamated free product of any countably many copies of $M$ over NN is Connes-embeddable. Then the amalgamated free generalized wreath product $M\wr^{\ast_{N}}_{\alpha}H$ is Connes-embeddable. In particular, if $H$ is a hyperlinear group, $\alpha:H\curvearrowright X$ is a sofic action, and $M$ is a Connes-embeddable tracial von Neumann algebra, then the free generalized wreath product $M\wr^{\ast}_{\alpha}H$ is Connes-embeddable, and under the same conditions, if we in addition have an amenable subalgebra NN of $M$, then the amalgamated free generalized wreath product $M\wr^{\ast_{N}}_{\alpha}H$ is Connes-embeddable.
@@
Let $H$ be a hyperlinear group, $\alpha:H\curvearrowright X$ be a sofic action, $A\leq G$ be an inclusion of countable discrete groups s.t. the amalgamated free product of any countably many copies of $G$ over AA is hyperlinear. Then the amalgamated free generalized wreath product $G\wr^{\ast_{A}}_{\alpha}H$ is hyperlinear. In particular, if $G,H$ are hyperlinear groups and $\alpha:H\curvearrowright X$ is a sofic action, then the free generalized wreath product $G\wr^{\ast}_{\alpha}H$ is hyperlinear, and under the same conditions, if we in addition have an amenable subgroup AA of $G$, then the amalgamated free generalized wreath product $G\wr^{\ast_{A}}_{\alpha}H$ is hyperlinear.